### v2

# Metastability and Ostwald Step Rule in the Crystallisation of Diamond and Graphite from Molten Carbon

Abstract
Experimental challenges in determining the phase diagram of carbon at temperatures and pressures near the graphite-diamond-liquid triple point are often related to the persistence of metastable crystalline or glassy phases, superheated crystals, or supercooled liquids. A deeper understanding of the crystallisation kinetics of diamond and graphite is crucial for effectively interpreting the outcomes of these experiments. Here, we reveal the microscopic mechanisms of diamond and graphite nucleation from liquid carbon through molecular simulations with first-principles machine learning potentials. Our simulations accurately reproduce the experimental phase diagram of carbon near the triple point and show that liquid carbon crystallises spontaneously upon cooling. Metastable graphite crystallises in the domain of diamond thermodynamic stability at pressures above the triple point. Furthermore, whereas diamond crystallises through a classical nucleation pathway, graphite follows a two-step process in which low-density fluctuations forego ordering. Calculations of the nucleation rates of the two competing phases confirm this result and reveal a manifestation of Ostwald’s step rule, where the strong metastability of graphite hinders the transformation to the stable diamond phase. Our results provide a new key to interpreting melting and recrystallisation experiments and shed light on nucleation kinetics in polymorphic materials with deep metastable states.
Introduction
The crystallisation of carbon from the melt under extreme conditions is highly relevant to Earth and planetary science,Timmerman et al. (2023); Ross (1981), materials manufacturing, and nuclear fusion research.Zylstra et al. (2022) The thermodynamic conditions near the graphite-diamond-liquid (GDL) triple point are especially of interest for geological and technological applications, but high-pressure flash heating experiments aiming to resolve this region of the phase diagram of carbon exhibit large discrepancies.Bundy et al. (1996); Togaya (1997); Zhao et al. (2013); Kondratyev and Rakhel (2019); Liang et al. (2019); Yang et al. (2023) The diverse technological applications of carbon materials stem from the different types of covalent bonding with tetrahedral patterns in diamond and tetrahedral amorphous carbon on one hand, and hexagonal motifs in graphite or graphitic carbon on the other. Such differences in covalent bonding engender large metastability of both common crystalline phases, graphite and diamond. These chemical features dictate the phase diagram of carbon, which exhibits a GDL triple point at about 12 GPa and 4500 K. Whereas the diamond-in-the-sky hypothesisRoss (1981) and recent experiments on nuclear fusion by laser shock compressionZylstra et al. (2022) have prompted several studies of the phase diagram of carbon and hydrocarbons at extreme pressures, the GDL triple point region has not been thoroughly addressed either by experiments or simulations. Conversely, understanding the thermodynamics and kinetics of phase changes in this region is of paramount importance for materials and planetary sciences. For example, in geosciences, diamonds reveal tectonic processes in Earth-like planets and the deep-earth cycle of carbon and other light elements.Timmerman et al. (2023) In materials processing, the liquid phase is an intermediate in the laser synthesis of carbon materials, including artificial diamonds, nanodiamonds, and N-vacancy doped diamonds for quantum computing, and its crystallisation behaviour upon cooling determines the properties of the synthesis products.Hull, Raj, and Saykally (2020)
Ostwald hypothesized that nucleation from the melt does not necessarily proceed directly to the thermodynamically most stable phase but through the phase that is separated by the lowest free energy barrier from the liquid.Ostwald (1897) This effect significantly accelerates crystallisation rates in the presence of metastable critical points.Talanquer and Oxtoby (1998) Molecular simulations of nucleation from the melt confirmed Ostwald’s step rule hypothesis in Lennard-Jones liquids, colloidal particles, and protein crystallisation.ten Wolde, Ruiz-Montero, and Frenkel (1995); Schilling et al. (2010); ten Wolde and Frenkel (1999) In the latter case, a two-step process was observed experimentally using dynamic light scattering and time-resolved spectroscopy.Vorontsova, Maes, and Vekilov (2015) Non-classical multi-stage nucleation pathways have also been discussed in the context of mineralisation from solution, with abundant numerical and experimental evidence,Gebauer et al. (2014), and in the case of polymorphic materials, the precipitation of metastable phases has been observed.Navrotsky (2004) Nevertheless, it is unexpected to encounter such a rich phenomenology in the homogeneous nucleation of a simple monoatomic system like carbon.
In this work, we investigate the kinetic aspects of diamond and graphite crystallisation, probing homogeneous nucleation from molten carbon at moderate pressures using machine-learning-accelerated, quantum-accurate molecular dynamics simulations. The range of pressures considered, between 5 and 30 GPa, corresponds to the pressure of the Earth’s mantle between 120 and 750 Km. Our simulations show the unexpected spontaneous crystallisation of metastable graphite at pressures up to 15 GPa, well within the domain of thermodynamic stability of diamond, and unravel the fundamentally distinct molecular mechanisms that lead to the nucleation of diamond and graphite. The crystallisation pathways exhibit a combination of these effects that are the consequence of Ostwald’s step rule and the large free energy barrier between graphite and diamond. These observations open new perspectives in the theory of crystallisation of single-component systems with metastable phases.
Results
Phase diagram and spontaneous crystallisation
We first discuss the phase diagram of carbon in the region of the GDL triple point (Figure 1a). Experimentally, the triple point was initially determined at T$=4100\pm 200$ K, and P$=12.5\pm 1$ GPa by extrapolating the graphite-diamond coexistence line to high temperatures.Kennedy and Kennedy (1976) Extrapolating the melting curves of graphiteTogaya (1997) and diamondBundy et al. (1996) suggests that the triple point should be at a higher temperature, between 4400 and 4800 K. However, in general, the experimental determination of the GDL triple point entails large uncertainties.Day (2012) In contrast to earlier measurements,Bundy (1963a) experiments now agree that the melting line of diamond has a positive slope while that of graphite has a negative slope at the triple point. The melting temperature of graphite also exhibits a maximum at $\sim$5 GPa. This was initially thought to be indicative of a liquid-liquid phase transition,Glosli and Ree (1999) which was not confirmed by further simulations or experiments.Wu et al. (2002); Wang, Scandolo, and Car (2005) Recent experiments suggest that the melting point of graphite at 2 GPa and that of diamond at 15 GPa may be over 6000 K, significantly higher than in previous works, adding further uncertainty to the carbon phase diagram and the location of the GDL triple point.Kondratyev and Rakhel (2019); Liang et al. (2019)
Here we compute the graphite-liquid and diamond-liquid coexistence lines, along with their metastable extensions, by large-scale phase-coexistence MD simulations using feed-forward neural network potentials fitted with an evolutionary strategy (neuroevolution potential, NEP)Fan et al. (2021) and trained on energies and forces computed at the level of density functional theory (DFT) either in the local density approximation (LDA)Deringer and Csányi (2017) and with a generalized gradient approximation functional with non-local van der Waals correlation ‘OptB88-vdW’.Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020) We dub these two models NEP@LDA and NEP@OptB88-vdW.
Our simulations closely reproduce the most commonly accepted experimental phase diagram regardless of the chosen NEP model.Bundy et al. (1996); Savvatimskiy (2015) In the NEP@LDA diagram, the intersection between the two melting curves establishes a GDL triple point at P=10.5$\pm$0.5 GPa and T=4650$\pm$50 K, at a slightly lower pressure than in most experiments but well within the temperature range. Starting from the triple point, we computed the graphite-diamond coexistence line using the Gibbs-Duhem integration method.Kofke (1993) The diamond-liquid coexistence line agrees with former FPMD simulations.Correa, Bonev, and Galli (2006); Wang, Scandolo, and Car (2005) Remarkably, our model reproduces the reentrant melting line of graphite in excellent agreement with measurements.Togaya (1997) This, to our knowledge, was never replicated in former molecular simulations. The reason for the negative slope of the melting line of graphite near the triple point is that the density of liquid carbon is higher than that of graphite at pressures higher than 2 GPa (Figure 1b). The density crossover at 2 GPa corresponds to the change of slope of the graphite melting line from positive to negative. This characteristic is extremely consequential for the crystallisation behaviour of graphite and diamond and is rarely captured by either empirical potentials or machine learning models,Ghiringhelli et al. (2008) with few notable exceptions, such as the revised long-range carbon bond-order potential (LCBOPII) and the environment-dependent interatomic potential (EDIP).Colonna et al. (2009); Marks (2000); Marchant et al. (2023) Notably, using different parameterisations of NEP does not have a significant effect on the melting lines and the location of the triple point. The phase diagrams reported in Figure S1 share the same features as those in Figure 1.
Since the approximations in the LDA exchange and correlation functional entail inaccuracies in the equations of state and the phase diagram of materials, usually overestimating the density of both solids and liquids,French and Mattsson (2014) we repeated the calculation of the phase diagram of carbon with NEP@OptB88-vdW.Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020); Fan et al. (2024) The NEP@LDA and NEP@OptB88-vdW phase diagrams are reported side by side in Figure S2. The GDL triple point with the NEP@OptB88-vdW model is at the same temperature as NEP@LDA but at a higher pressure (P=17 GPa and T=4660 K). In general, the two phase diagrams exhibit the same features with just a pressure offset of $\sim 6.5$ GPa.
As expected from previous DFT calculations,French and Mattsson (2014) the equations of state of liquid carbon, computed at 5000 and 7000 K differ with NEP@LDA overestimating and NEP@OptB88-vdW underestimating the density compared to experiments (Figure S3).Kondratyev and Rakhel (2019) However, the structure of the liquid, characterised through radial distribution functions (Figure S4) and the analysis of the local coordination (Figure S5), exhibits the same behaviour as the pressure is increased. At any pressure between 5 and 30 GPa, compared to the NEP-OptB88-vdW liquid, the NEP@LDA liquid exhibits a larger number of tetrahedrally coordinated carbons that increases linearly with the pressure, and a lower number of twofold coordinated carbons that decreases. The number of threefold coordinated carbons has a maximum between 5 and 10 GPa for NEP@LDA and above 15 GPa for NEP@OptB88-vdW. The differences in the properties of the liquids obtained with the two models may be reconciled with a pressure shift of about 10 GPa. The liquid structure is consistent with previous studies based on DFT or accurate empirical potentials.Ghiringhelli et al. (2005, 2008); Colonna et al. (2009)
Due to the small size of molecular models and the use of periodic boundary conditions, MD simulations customarily overestimate the stability of liquid phases below the melting point and the formation of amorphous phases by quenching from the melt. However, MD runs in which liquid carbon models equilibrated at 5000 K are quenched at the rate of 60 K/ns exhibit spontaneous crystallisation into either diamond or graphite, depending on the pressure. The temperatures at which spontaneous crystallisation was observed are reported in the phase diagram (Figure 1a and S2) and define the lowest-temperature boundary of the region of metastability of supercooled liquid carbon. This leads us to conclude that liquid carbon, as opposed to other group IV elemental liquids like silicon and germanium, is not a good glass former at high pressure. Therefore, quenching from the melt at high pressure is not a viable route to obtain tetrahedral amorphous carbon. The temperature of spontaneous crystallisation of graphite decreases with pressure with a more accentuated negative slope than the phase coexistence line. Conversely, the temperature of spontaneous crystallisation of diamond increases with pressure but with a similar slope as the melting curve.
Strikingly, we observe spontaneous crystallisation of graphite at pressures as high as 15 GPa, which is about 5 GPa higher than the pressure of the GDL triple point and 7 GPa higher than the graphite-diamond phase boundary at the crystallisation temperature. The occurrence of a metastable structure in the crystallisation pathway is predicted by Ostwald’s step rule when the metastable phase has a lower free energy barrier than the thermodynamically stable one.Ostwald (1897); Stranski and Totomanow (1933) This condition corresponds to a closer structural similarity between the nucleating metastable structure and the parent liquid. We observe that the highest pressure at which graphite crystallizes spontaneously corresponds to the condition where the density of the liquid is midway between that of graphite and that of diamond (see Figure 1b). This observation suggests that the density of the liquid is an important driver of crystallisation. In turn, the change of preference for spontaneous crystallisation from graphite to diamond is not dictated by any abrupt changes in the structure of liquid carbon: the equation of state, pair correlation function, and local coordination undergo gradual changes as the pressure increases (Figures S3-S5) in accordance with previous studies.Ghiringhelli et al. (2005); Colonna et al. (2009) It is worth noting that at these pressures, the liquid is still predominantly three-fold coordinated. In most cases, the metastable nucleus transforms spontaneously into the stable phase as it grows. In our simulations, however, graphite persists, as the activation free energy to transform graphite into diamond is high, and crystallisation happens too rapidly. The crystallisation of metastable crystalline phases from rapidly quenched liquids has been observed for minerals and alloys, especially when simple low-density structures compete with complex structures, but it was not formerly hypothesised for carbon. The pressure conditions between the graphite-diamond phase coexistence line and 15 GPa are insufficient to overcome the high barrier to transforming graphite into diamond over the typical MD timescales.Khaliullin et al. (2011); Zhu et al. (2020) However, experimentally graphite transforms directly into diamond upon compression at 12.5 GPa above 3000 K,Bundy (1963b) thus suggesting that graphite would be a long-lived metastable intermediate state in the crystallisation of diamond.
Spontaneous crystallisation of metastable graphite above the graphite-diamond coexistence line also occurs in simulations performed with the NEP@OptB88-vdW model (Figure S2). In this case, we observed metastable nucleation of graphite up to 19 GPa at 3670 K, where the corresponding graphite-diamond phase boundary is at 14.5 GPa (Figure S2). Since the two NEP models provide the same qualitative physical picture of the thermodynamics of carbon in the region of interest and of the kinetics of crystallisation, we restrict the analysis in the following sections to results obtained using the NEP@LDA model.
Nucleation rates
To validate the observed metastable crystallisation of graphite from liquid carbon, we computed the nucleation rates of graphite and diamond at moderate supercooling. These calculations require the use of an enhanced sampling technique with the possibility of biasing the nucleation of either phase. For this purpose, we used forward flux sampling molecular dynamics.Li et al. (2009); Li, Donadio, and Galli (2009) To identify the graphite and diamond nucleating crystallites, we adopted local order parameters based on the spherical harmonic expansion of the atomic environment ‘$q_{l}$-dot".ten Wolde, Ruiz-Montero, and Frenkel (1996); Ghiringhelli et al. (2007) For diamond, we use $l=6$ and for graphite $l=3$ with the additional constraint that atoms must be three-fold coordinated. Figure 2(d-f) shows the nucleation rates of graphite and diamond as a function of temperature up to 4100 K at $P=$13, 15, and 17 GPa. These calculations confirm a preference for graphite crystallisation at any temperature at 13 GPa. Nucleation rates for the two phases are similar at 15 GPa with a slight preference for graphite, whereas crystallisation of diamond is predominant at 17 GPa.
The rates and the estimates of the critical size of the nuclei obtained by FFS allow us to fit a model based on classical nucleation theory (CNT). In CNT, the free energy of formation of a crystalline nucleus of size $N$ in a liquid is the result of a negative term that accounts for the difference in chemical potential between the solid and the liquid ($\Delta\mu$) and a positive term that accounts for the cost of creating solid/liquid interface. The CNT rate can be expressed as:
$R_{CNT}(T)=A\exp{\left(-\frac{\Delta G^{\star}}{k_{B}T}\right)}$ (1)
where $A$ is a kinetic prefactor and $\Delta G^{\star}$ is the free energy of formation of the critical nucleus:
$\Delta G^{\star}=\frac{16\pi\gamma_{LS}^{3}}{3\rho^{2}\Delta\mu^{2}}$ (2)
In this expression, $\rho$ is the number density of the solid, and $\gamma_{LS}$ is the average interfacial tension between the liquid and the solid phase. $\Delta\mu$ is usually approximated as $\Delta\mu=\Delta H_{m}(T_{m}-T)/T$ where the enthalpy of melting ($\Delta H_{m}$) and the melting temperature ($T_{m}$) are calculated from MD simulations at solid/liquid coexistence conditions, leaving $A$ and $\gamma_{LS}$ as the only unknowns. Figure 2(d-f) shows that CNT (continuous lines) fits the FFS nucleation rates of both graphite and diamond. We note that the fitting of the FFS rate against CNT does not rely on the assumption of a spherical critical nucleus, but only requires the shape to remain unchanged at different temperatures. $\Delta G^{\star}$ obtained from this fit provides an estimate of the effective liquid/solid interfacial tension for graphite and diamond nuclei assumed spherical, where the former is $\gamma_{LS,G}=1.15$ J/m2 and the latter is $\gamma_{LS,D}\sim 2.0$ J/m2 at $P=15$ GPa. At an atomistic level, the high value of $\gamma_{LS,D}$ at moderate pressure can be explained by the stark difference in the bonding environment between the liquid, mostly three-fold coordinated, and the four-fold coordinated diamond crystallite.Ghiringhelli et al. (2007) The much smaller value of $\gamma_{LS,G}$ suggests that the direct nucleation of diamond can occur only when the chemical potential term is dominant, i.e., when the liquid is substantially more supercooled with respect to diamond than to graphite.
[…]
Nucleation pathways
To study unbiased nucleation pathways and obtain an estimate of the nucleation rates of graphite and diamond at deep supercooling,Yuhara et al. (2015) we have performed direct MD simulations at pressures of 15 GPa and 15.5 GPa and a temperature of 3650 K. Averaging over 10 runs at each pressure, we computed a mean first passage time (MFPT) $\tau_{MFPT}=21.1\pm 2.8$ ns corresponding to a nucleation rate of 1.7±$\pm 0.2\times 10^{33}$ s-1m-3 for graphite at 15 GPa, and $\tau_{MFPT}=16.3\pm 5.4$ ns corresponding to 2.3±$\pm 0.7\times 10^{33}$ s-1m-3 for diamond at 15.5 GPa. The (MFPT) nucleation rate of graphite is found to match well the CNT model fitted through the FFS rates (Figure 2e), further confirming the validity of CNT and the applicability of FFS for graphite nucleation.
Comparing the qualitative features of the growing nuclei in direct simulations to FFS confirms that the chosen order parameters to define the flux interfaces do not bias the sampling towards unrealistic nucleation pathways. Figure 3 shows snapshots along the crystallisation pathway of graphite and diamond from unbiased simulations in which liquid carbon is cooled at 15 and 15.5 GPa. Diamond nucleates into compact crystallites with an isotropic aspect ratio, which justifies the spherical nucleus approximation often adopted in CNT. Similar behaviour was observed in the crystallisation of other tetrahedral systems, including silicon, germanium, and water ice.Li, Donadio, and Galli (2009); Li et al. (2011) This also means that the average estimate of $\gamma_{LS,D}$ is reliable. Conversely, graphite nucleation proceeds through the formation and stacking of small graphite patches which produce anisotropic crystallites elongated along the cross-plane axis of graphite (Figure 3(top). The growing graphite crystals are highly non-spherical, and the free energy of the interface between the different facets and the liquid dictates their shape.
[…]
Prenucleation mechanism
Previous considerations indicate that density is the main similarity parameter that drives the crystallisation of graphite or diamond from molten carbon, since both transitions involve large density variations. At the tipping pressure of 15 GPa, the crystallisation of graphite is accompanied by a density drop from 2.9 to 2.5 g/cm3, whereas when diamond is formed, the density rises to 3.5 g/cm3. Such extreme density differences between the liquid and the solid phases affect nucleation pathways. Previous studies showed that, following Ostwald’s step rule, nucleation could be a two-step process in which density fluctuations with the formation of high-density amorphous or liquid aggregates precede crystalline ordering. This effect was observed during crystal nucleation both in supercooled single-component liquids and supersaturated solutions.ten Wolde and Frenkel (1997, 1999); Schilling et al. (2010); Hu and Tanaka (2022); Finney and Salvalaglio (2022) We observe a similar effect in the crystallisation of graphite. Figure 4(a) shows the probability that low-density clusters and crystalline nuclei of a given size form spontaneously in the liquid at 3650 K and 15 GPa. The shape of the distribution indicates that the crystallinity and low-density order parameters do not correlate. The occurrence of crystalline clusters without local low-density fluctuations can be explained by the fact that the system tends to form diamond-like small nuclei. However, spontaneous nucleation pathways to graphite, one of which is represented in the graph as a red dotted line, evolve through the formation of a low-density cluster that eventually develops into a crystalline nucleus. In the case of diamond nucleation, instead, as shown in Figure 4(b), the density fluctuations producing high-density clusters are correlated to the crystallinity order parameter. Consequently, the crystalline nucleus is not preceded by a higher-density disorder precursor: local densification and ordering coincide. This difference further supports the evidence that graphite nucleation follows a lower-barrier free energy pathway, facilitated by local low-density fluctuations in liquid carbon even at the thermodynamic conditions at which diamond is the stable phase.
Discussion
Accurate and efficient machine learning potentials trained on density functional theory calculations, enabling a total of several microseconds of MD simulations of systems of several thousand atoms, provided us with an invaluable tool to explore the phase diagram of carbon near the GDL triple point and the intricate kinetics of homogeneous nucleation of graphite and diamond from the melt. Whereas liquid carbon is often considered a glass former, we observed spontaneous crystallisation of both graphite and diamond in unbiased MD simulations in which the liquid is cooled rapidly at constant pressure. This suggested that liquid carbon at pressures between 5 and 30 GPa is not a glass former, and provided us with unbiased trajectories to examine the nucleation pathways at the molecular scale. We observed an unexpected wealth of physical phenomena related to homogeneous nucleation in a deceivingly simple monoatomic system such as liquid carbon. These phenomena are inherently connected to Ostwald’s step rule, leading to highly non-classical nucleation effects. In MD simulations, graphite crystallises spontaneously in the domain of thermodynamic stability of diamond due to the smaller density differences between liquid carbon and graphite and the lower liquid/solid interfacial free energy. Depending on the details of the potential, metastable graphite may crystallize up to a pressure between 5 and 7.5 GPa above the graphite-diamond coexistence line. In this range of pressures, graphite is a long-lived metastable state on the way to the crystallisation of diamond.Bundy (1963b) Additionally, we found graphite nucleates through a non-classical two-step process where crystalline ordering is preceded by the formation of a low-density liquid region that further lowers the barrier to nucleation. Large differences in the interfacial free energy of different facets dictate the strongly anisotropic shape of growing graphite crystallites. Conversely, at higher pressure, diamond nucleates through a classical one-step process, forming densely packed tetrahedrally bonded crystalline spherical clusters. The two intrinsically different nucleation mechanisms of graphite and diamond are schematically represented in Figure 4c. Even if graphite nucleation is non-classical, a CNT fit of the nucleation rates yields physically meaningful parameters for both graphite and diamond, suggesting that even non-classical nucleation processes can be mapped on a CNT model.Zahn (2015)
The insight into homogeneous nucleation from carbon melt provided by our simulations may help interpret inconsistencies among historical electrical and laser flash-heating experiments aimed at resolving the phase diagram of carbon near the GDL triple point. Depending on the details of these experiments and the recrystallisation conditions, the system may remain trapped in metastable graphitic configurations. These observations may also impact the manufacturing of carbon-based materials such as synthetic diamonds and nanodiamonds at high pressure and high temperature.
Methods
.1 Machine learning potentials
Molecular dynamics simulations were performed using the GPUMD v3.6 code with a NEP3 neuroevolution potentialFan et al. (2021, 2022) fitted to density functional theory (DFT) calculations in the local density approximation (LDA) for the exchange and correlation functional (dubbed NEP@LDA) and a NEP4 potential developed based on the OptB88-vdW functional (NEP@OptB88-vdW).Klimeš, Bowler, and Michaelides (2010); Rowe et al. (2020) For both potentials, the databases of configurations, energies, forces, and virials were formerly generated to study diverse carbon systems, including graphite, diamond, low and high-density carbon, and liquid carbon.Deringer and Csányi (2017); Rowe et al. (2020) We conducted extensive tests to verify that the resulting NEP potentials are suitable for modelling diamond and graphite. Besides the phase diagrams shown in Figure 1 and S2, we verified that the NEP@LDA reproduces the structural, vibrational, and elastic properties of diamond and graphite accurately (Table S1). These results suggest that, whereas LDA is the lowest-level approximation for exchange and correlation in DFT, it accurately describes the mechanical and thermodynamic properties of graphite, particularly the liquid-graphite phase coexistence. In the pressure range of interest, i.e., above 10 GPa, our predictions for the melting line of diamond are within 100 K of DFT calculations with GGA functionals.Wang, Scandolo, and Car (2005); Correa, Bonev, and Galli (2006)
Additionally, we verified the sensitivity of our predicted phase diagram and spontaneous crystallisation conditions upon the hyperparameters and the training process of the NEP@LDA potentials. We generated three more carbon NEP@LDA potentials varying several hyperparameters, including cutoff (cutoff radii for the radial and angular descriptor components), n_max (number of radial functions for the radial and angular descriptor components), basis_size (number of radial basis functions for the radial and angular descriptor components), population size, and number of generation. Figure S1 shows that the melting line of graphite may shift by at most 100 K, whereas the melting line of diamond is insensitive to the different NEP parameterisations. Accordingly, crystallisation temperatures undergo variations up to 150 K, including the uncertainty inherent to the stochastic nature of this process. The highest pressure at which metastable graphite crystallises spontaneously varies between 13 and 15 GPa. The hyperparameters for training the NEP@LDA and NEP@OptB88 models are reported in Table S2.
.2 Molecular Dynamics
MD simulations probing the properties of liquid carbon and spontaneous crystallisation were performed using the GPUMD package for cubic systems containing 4096 carbon atoms.Fan et al. (2022) The equations of motion were integrated with a timestep of 0.5 fs. We have verified that with a 0.5 fs timestep, MD simulations of liquid carbon at $\sim 4500$ K conserve the energy within 1 meV/atom over 100 ps (Figure S6). The isobaric canonical ensemble (constant number of particles, pressure, and temperature - NPT) was controlled using the stochastic rescaling scheme with coupling times of 1 ps for the temperature and 5 ps for the pressure.Bussi, Donadio, and Parrinello (2007); Bernetti and Bussi (2020) Spontaneous crystallisation of diamond and graphite was observed in direct MD simulations of liquid carbon at constant pressure, in which the temperature was ramped from 5,000 K to 3,500 K in 25 ns, corresponding to a cooling rate of 60 K/ns. These simulations were initiated from liquid models equilibrated at 5000 K for at least 5 ns. The structural features of these liquids are characterised by the radial distribution functions and the coordination numbers reported in Figures S4 and S5.
.3 Calculation of the phase boundaries
[…]
To compute the graphite-diamond phase boundary and to refine the solid-liquid coexistence lines, we integrated the Gibbs-Duhem equationKofke (1993) by finite differences in $\beta$:
$\frac{dp}{d\beta}=-\frac{\Delta h}{\beta\Delta v}$ (3)
where $\beta=1/k_{B}T$, $\Delta h$ is the difference in molar enthalpy between the two phases, and $\delta v$ is the difference in molar volume. For the graphite-liquid coexistence line, we integrated Eq. 3 starting from the triple point with a fixed $\Delta\beta=0.05$ eV-1. At each step, $\Delta h$ and $\Delta v$ are estimated as ensemble averages over the second half of 100 ps NPT runs.
.4 Nucleation rates from forward flux sampling and order parameters
Nucleation rates of diamond and graphite are computed using our implementation of forward flux sampling (FFS) in LAMMPS, based on carbon NEP. In FFS, nucleation rate $R$ is given as the product of $\Phi_{\lambda_{0}}$, which is the initial flux rate of nucleation trajectory crossing the initial milestone $\lambda_{0}$, and $P(\lambda_{F}|\lambda_{0})$, which measures the probability for a trajectory starting from $\lambda_{0}$ and successfully reaching the final milestone $\lambda_{F}$. $\Phi_{\lambda_{0}}$ is obtained by $N_{0}/t_{0}V$, where $N_{0}$ (120) is the number of successful crossings collected at $\lambda_{0}$, $V$ is the volume of the simulation cell, $t_{0}$ is the total time for $N_{0}$ crossings. $P(\lambda_{F}|\lambda_{0})=\prod_{i=1}^{n}P(\lambda_{i}|\lambda_{i-1})$, where $P(\lambda_{i}|\lambda_{i-1})=N_{i}/M_{i-1}$ is computed through conducting $M_{i-1}$ MD trial runs starting from the milestone $\lambda_{i-1}$ and collecting $N_{i}$ (120) successful crossing at the next adjacent interface $\lambda_{i}$. For each thermodynamic condition $T/P$, three independent FFS runs are carried out to obtain the geometric mean rate $\langle R\rangle=(\prod_{i=1}^{3}R_{i})^{1/3}$ and the standard error of $\ln\langle R\rangle$ is given by $\sqrt{\sum_{i=1}^{3}(\ln R_{i}-\ln\langle R\rangle)^{2}}/3$.Zhao and Li (2023)
In FFS, the milestones $\lambda_{i}$ are defined based on the size of the largest crystalline cluster in the melt. Three types of carbon atoms, namely, liquid-like, diamond-like, and graphite-like, are differentiated through a local bond-order parameter $q_{n}$.Li, Donadio, and Galli (2009) Similar to the Steinhardt bond order parameter,Steinhardt, Nelson, and Ronchetti (1983) $q_{n}$ is rotationally invariant; however, as a local order parameter, $q_{n}$ reflects the inherently local character of nucleation. A diamond-like carbon is defined as the carbon atom that has a $q_{6}>0.5$ (see Figure S8) and exactly 4 nearest neighbours, whereas a graphite-like carbon is the one that has a $q_{3}<-0.85$ (see Figure S8) and exactly 3 nearest neighbours. These choices are made based on both the distributions of $q_{n}$ for the three phases and the characteristic local coordination in diamond and graphite. For the $q_{n}$ analysis, the cut-off distance in identifying the nearest neighbours in both diamond and graphite is set to be 1.8 Å. In determining the size of a crystalline cluster through the connectivity analysis, the same cut-off distance 1.8 Å is used for diamond, whereas a greater cut-off distance 3.5 Å is used for graphite, to account for the inter-layer spacing between neighbouring graphene sheets.
.5 Interfacial tension calculations
Due to the thermal fluctuations, the interface between distinct phases is not entirely flat. According to capillary wave theory, these fluctuations can be connected to the interfacial tension in the following manner:
$\mathrm{\gamma}\equiv\frac{k_{B}T}{2\pi\langle\sigma^{2}\rangle}\ln{\frac{L}{\xi}},$ (4)
where $T$ is the absolute temperature, $k_{B}$ is the Boltzmann constant, $\sigma^{2}$ is the mean squared fluctuation of molecules at the interface, $L$ is determined by the size along the $x$- or $y$- dimension (assuming $z$ is normal to the surface) and $\xi$ is the bulk correlation length.Cordeiro (2003) Using this relationship, we are able to estimate the interfacial free energy differences between the liquid-graphite (basal facet) and the liquid-graphite (prismatic facet), using the phase coexistence MD simulations used to compute the liquid-solid phase boundary. We set $L=38.046$ according to our simulation cell size, and $\xi=1.9$ corresponding to the first minimum of the radial distribution function of liquid carbon. $\sigma^{2}$ is calculated as: $\langle\sigma^{2}\rangle\equiv\langle(z-\langle z\rangle)^{2}\rangle$ where we selected atoms from the largest liquid carbon cluster belonging to one of the two interfaces in the periodic cell. The calculation is halted when the chosen interface becomes locally vertical (i.e., the approximate tangent plane contains $\hat{z}$) or makes contact with the other interface. We carried out the comparison between the two facets for three simulations, each held at 10 GPa with varying temperatures (4600 K, 4700 K, and 4800 K).
Data Availability Statement
The data that support the findings of this study are available open-source at http://doi.org/10.5281/zenodo.15567723. The implementation of forward flux sampling in LAMMPS is available at https://github.com/WanyuZhao/FFS-LAMMPS. Source data are provided with this paper.
Acknowledgements.
We are grateful to Gabriele C. Sosso for providing a critical assessment of the manuscript. This work was supported by the National Science Foundation under Grants No. 2053235 (DD) and 2053330 (TL).
Author Contributions Statement
DD and TL conceived the study. DD and WZ conducted the molecular dynamics simulations. SC trained the machine learning models. WZ, MLB, SC, DD, and TL analysed the results. DD drafted the manuscript. All authors contributed to editing and finalizing the manuscript.
Competing Interests Statement
All authors declare no competing interests.